# Simple Methods Defend RAG Systems Well Against Real-World Attacks

### v1

@@ Appendix A Prompts
### Instructions: 1. **Review the Knowledge Bases**: Carefully examine the knowledge base and adversarial knowledge base provided below. Identify key topics such as COVID-19, vaccines, public health measures, and related scientific inquiries. 2. **Analyze the Question**: Evaluate the given question for its relevance to the knowledge base. Consider both specific keywords and the overall context and intent. Identify any implicit context or background information that might be relevant. 3. **Determine Domain Classification**: - **In-Domain**: The question can be answered using the knowledge base. It must directly relate to the knowledge baseś focus areas, such as COVID-19 vaccines, public health measures, and related scientific inquiries. - **Out-of-Domain**: The question falls outside the scope of the knowledge base. This includes questions requiring speculative predictions, future events, or information not covered by the static knowledge base. 4. **Explicit Decision Criteria**: Define clear criteria for classifying questions as ”In-Domain” or ”Out-of-Domain.” Include specific keywords, topics, or themes that are considered in-domain, and provide examples of out-of-domain questions. For example, questions about speculative future events or unrelated scientific fields are out-of-domain. 5. **Provide a Clear Explanation**: Offer a detailed explanation for your decision, referencing specific examples or sections from the knowledge base when applicable. Use evidence from the knowledge base or other authoritative sources to support your decision. Reference specific studies, guidelines, or expert opinions to enhance specificity. 6. **Handle Ambiguities**: Recognize and address potential ambiguities in inputs. Hypothesize potential meanings for ambiguous terms and evaluate their relevance to the domain. Clarify reasoning based on the specific context of the task. 7. **Sensitivity and Bias Awareness**: Approach sensitive terms, especially those related to identity, ethnicity, or religion, with care. Ensure explanations do not perpetuate bias or insensitivity. 8. **Example Integration**: Use examples of both in-domain and out-of-domain inputs to guide your reasoning and improve differentiation between relevant and irrelevant inputs. Consider potential counterexamples or scenarios where a term might be out-of-domain. 9. **Logical Structure Guidance**: Follow a structured reasoning framework. Start with identifying domain criteria, analyze the input against these criteria, consider counterexamples, and conclude with a well-supported decision. 10. **Explanation Depth**: Provide detailed explanations for decisions. Discuss the absence of connections to the knowledge base topics when classifying an input as out-of-domain. Acknowledge the inherent uncertainty in certain types of questions. 11. **Feedback Loop**: Implement a feedback loop to learn from past decisions and adjust your reasoning process to avoid repeating mistakes. Reflect on your reasoning and consider feedback for continuous improvement in future evaluations. 12. **Keyword and Phrase Analysis**: Perform a detailed analysis of keywords and phrases. Compare these with the input question to justify classification decisions based on the presence or absence of domain-specific keywords. 13. **Consideration of Counterarguments**: Consider alternative perspectives or counterarguments that might suggest a question is out-of-domain. Address these counterarguments to provide a balanced and comprehensive analysis.
### Response Format: - **Decision**: [In-Domain/Out-of-Domain] - **Explanation**: [Provide a clear reason based on the question and knowledge base.]
@@
Figure 16: This prompt is designed to force GPT-4o to act an out-of-domain detector. Optimized by (Yuksekgonul et al. 2025).
@@ Appendix C Further Results
Further Domains Results
@@
EVR criterion p-values criterion Dataset $\epsilon$-ball $\epsilon$-cube $\epsilon$-rect LogReg SVM GMM $\epsilon$-ball $\epsilon$-cube $\epsilon$-rect LogReg SVM GMM COVID-19 0.942 0.937 0.85 0.981 0.985 0.937 0.985 0.951 0.942 0.966 0.976 0.942 SU 0.94 0.928 0.936 0.967 0.972 0.964 0.944 0.944 0.904 0.964 0.976 0.96 history 0.899 0.894 0.893 0.963 0.962 0.928 0.933 0.932 0.775 0.966 0.965 0.951 crypto 0.913 0.912 0.896 0.966 0.967 0.976 0.908 0.897 0.772 0.968 0.969 0.981 chess 0.916 0.906 0.867 0.959 0.96 0.948 0.923 0.917 0.855 0.961 0.961 0.959 cooking 0.916 0.91 0.894 0.983 0.982 0.977 0.922 0.908 0.778 0.984 0.985 0.98 astronomy 0.933 0.928 0.918 0.975 0.974 0.977 0.922 0.93 0.834 0.975 0.975 0.971 fitness 0.941 0.933 0.898 0.967 0.967 0.95 0.936 0.94 0.767 0.97 0.97 0.962 anime 0.866 0.85 0.838 0.939 0.941 0.878 0.844 0.861 0.815 0.932 0.942 0.889 literature 0.898 0.901 0.874 0.953 0.946 0.905 0.913 0.898 0.888 0.929 0.933 0.895 biomedical 0.915 0.908 0.879 0.93 0.929 0.904 0.87 0.852 0.714 0.937 0.939 0.948 music 0.908 0.906 0.872 0.927 0.933 0.892 0.912 0.904 0.864 0.941 0.944 0.917 film 0.896 0.896 0.896 0.94 0.938 0.923 0.887 0.879 0.848 0.942 0.943 0.919 finance 0.905 0.898 0.869 0.938 0.94 0.88 0.884 0.885 0.803 0.944 0.946 0.935 law 0.869 0.863 0.851 0.903 0.922 0.877 0.849 0.832 0.78 0.908 0.916 0.908 computing 0.884 0.869 0.847 0.917 0.921 0.872 0.852 0.862 0.845 0.925 0.922 0.878 Average 0.909 0.902 0.88 0.951 0.952 0.924 0.901 0.900 0.824 0.951 0.954 0.937 Average #PCs 7 6 4 140 143 40 13 15 5 125 136 65
Table 6: The results for both EVR and p-values criteria in 16 domains. The results are similar with these of Table 1, but unaggregated.
Then, we show the detailed results for each combination of positive and negative samples in the training set and for different test sets across our three datasets: COVID-19 (always treated as positive), LLM-attack, and 4chan. The results are shown in Tables 7-14.
@@
For robustness, we experimented with different models as embedding extractors. More specifically, we employed ModernBERT with a pooling mechanism to extract sentence embeddings and NovaSearch/stella_en_400M_v5, which has shown state-of-the-art results in many NLP tasks. The experiments focused on the p-values criterion and the training sets of COVID-19 samples as positive and a mixture of all datasets as negative. The results are shown in Table 15,  16,  17 for the testing sets of a mixture of all datasets, 4chan, and LLM-Attack, correspondingly. As observed, the performance does not have significant fluctuations among the different models, indicating the effectiveness of a simplistic model that is used throughout our main findings.
@@
Further Relevance & Correctness Results
@@
Table 19: Mean ($\pm$±Standard Deviation) of both dimensions for the different groups of in-domain (ID) and out-of-domain (OOD) questions for the COVID-19 domain, annotated by humans.
@@
We conducted additional experiments focusing exclusively on Correctness to gain deeper insights into potential errors. Initially, we annotated all samples independently. Subsequently, we replaced the LLM-as-a-Judge with the Claude-3.5-Sonnet model. Interestingly, the new judge independently identified the necessity of a third category, ”N/A”, and classified 18 cases accordingly. To further validate this, we also allowed ourselves to mark ”N/A” cases independently of Claude. This resulted in a high agreement between our annotations and the model’s judgments, indicating 12 common cases as ”N/A”. The results of both approaches are presented in Table 22.
@@
Table 22: Results for Correctness with samples annotated entirely by us and by Claude. ”N/A” cases mean that these cases cannot be judged scientifically. 12 samples were noted as ”N/A” from both our and Claude annotation.
As shown in the prompts in Fig.6 and Fig.7, two different prompts were used for the Generator component. This approach was necessary because using a single prompt proved ineffective in maximizing the number of questions that could be answered. To address this, we first employed the prompt that yielded the highest number of responses (Fig.6). The unanswered queries from this round were then processed using a second prompt (Fig.7). In the initial round, 281 out of 300 queries were successfully answered. The remaining 19 queries were reattempted using the second prompt across 10 iterations, resulting in 16 additional responses. Table 23 illustrates the results when all 300 samples are utilized. Table 24 presents the results when only the initial 281 answered queries are considered. Table 25 reports the results for the 16 additional responses generated by the second prompt, and Table 26 provides the final results for all 297 answered queries. The three unanswered queries were met with the response: ”I can’t assist with this question.”
@@
Table 23: Mean ($\pm$±Standard Deviation) of both dimensions for the different groups of in-domain (ID) and out-of-domain (OOD) questions for COVID-19 domain, without excluding the ”N/A” cases.
@@
Table 24: Mean ($\pm$±Standard Deviation) of both dimensions for the different groups of in-domain (ID) and out-of-domain (OOD) questions, when only the 281 initially answered queries are considered.
@@
Table 25: Mean ($\pm$±Standard Deviation) of both dimensions for the different groups of in-domain (ID) and out-of-domain (OOD) questions, when only the 16 questions were answered after multiple rounds and using a new prompt.
@@
Table 26: Mean ($\pm$±Standard Deviation) of both dimensions for the different groups of in-domain (ID) and out-of-domain (OOD) questions, where the 297 questions are considered, excluding the cases that were responded with ”I can’t assist with that”.
@@
In Table 29, we present the complete results for all developed methods evaluated on the 300 samples from our second study. This table complements Table 4, which exclusively reports the results for the GMM as it achieved the best performance. Similarly, in Table 30, we provide the results for the additional variations of the GPT-4o evaluator that we experimented with—specifically, the setting where only positive samples were included in the prompt and the 20-shot learning approach. Finally, in Table 31, we present our prompt optimization results. For optimization purposes, we developed 3 additional prompts (Fig. 13 - 15). The last two (Fig. 14 and  15) make use of zero-shot CoT (Chain-of-Thought) (Kojima et al. 2022). For even further optimization, we utilize the methodology of (Yuksekgonul et al. 2025), which results in the prompt of Fig. 16. As we observe, our best results are achieved by the prompt of the main paper. In addition, we leveraged this best prompt and developed an extensively utilized method. Specifically, we ran five passes of our best prompt using a temperature of 0.7. We then applied a majority vote strategy across the five predictions to reduce uncertainty in the classification of in-domain (ID) vs. out-of-domain (OOD) queries. The updated result is included as a new row (“UA” – Uncertainty-Aware). We find that the UA setup correctly classifies one additional query in total compared to the Main Prompt setup, indicating a minor improvement. This suggests that while incorporating stochasticity and ensembling can add value, our proposed methods remain highly competitive, especially given their interpretability and computational efficiency.
@@
Table 32: Mean ($\pm$±Standard Deviation) of both dimensions for the different groups of in-domain (ID) and out-of-domain (OOD) questions for the SU domain, using GPT-4o as a generator. We notice that this regards only LLM-as-a-Judge evaluation.
@@ Appendix D Meaning Of PCs
This manual examination was followed by a verification step using GPT-4o, which confirmed the identified patterns. In Tables 33 and 34, we present the discovered patterns for the COVID-19 and SU domains, respectively. Each table includes: (1) the experimental condition in which the PC was selected, (2) a brief label describing the dominant pattern that the PC appears to capture, and (3) three example queries with the highest scores along that PC.

### v2

@@ Appendix A Prompts
You are an AI tasked with determining whether a given question is **in-domain** or **out-of-domain** relative to a provided knowledge base. The knowledge base consists of a list of questions that represent the domain of knowledge you are familiar with. You are also given an adversarial knowledge base containing questions designed to be out-of-domain and potentially harmful. ### Instructions: 1. **Review the Knowledge Bases**: Carefully examine the knowledge base and adversarial knowledge base provided below. Identify key topics such as COVID-19, vaccines, public health measures, and related scientific inquiries. 2. **Analyze the Question**: Evaluate the given question for its relevance to the knowledge base. Consider both specific keywords and the overall context and intent. Identify any implicit context or background information that might be relevant. 3. **Determine Domain Classification**: - **In-Domain**: The question can be answered using the knowledge base. It must directly relate to the knowledge baseś focus areas, such as COVID-19 vaccines, public health measures, and related scientific inquiries. - **Out-of-Domain**: The question falls outside the scope of the knowledge base. This includes questions requiring speculative predictions, future events, or information not covered by the static knowledge base. 4. **Explicit Decision Criteria**: Define clear criteria for classifying questions as "In-Domain" or "Out-of-Domain." Include specific keywords, topics, or themes that are considered in-domain, and provide examples of out-of-domain questions. For example, questions about speculative future events or unrelated scientific fields are out-of-domain. 5. **Provide a Clear Explanation**: Offer a detailed explanation for your decision, referencing specific examples or sections from the knowledge base when applicable. Use evidence from the knowledge base or other authoritative sources to support your decision. Reference specific studies, guidelines, or expert opinions to enhance specificity. 6. **Handle Ambiguities**: Recognize and address potential ambiguities in inputs. Hypothesize potential meanings for ambiguous terms and evaluate their relevance to the domain. Clarify reasoning based on the specific context of the task. 7. **Sensitivity and Bias Awareness**: Approach sensitive terms, especially those related to identity, ethnicity, or religion, with care. Ensure explanations do not perpetuate bias or insensitivity. 8. **Example Integration**: Use examples of both in-domain and out-of-domain inputs to guide your reasoning and improve differentiation between relevant and irrelevant inputs. Consider potential counterexamples or scenarios where a term might be out-of-domain. 9. **Logical Structure Guidance**: Follow a structured reasoning framework. Start with identifying domain criteria, analyze the input against these criteria, consider counterexamples, and conclude with a well-supported decision. 10. **Explanation Depth**: Provide detailed explanations for decisions. Discuss the absence of connections to the knowledge base topics when classifying an input as out-of-domain. Acknowledge the inherent uncertainty in certain types of questions. 11. **Feedback Loop**: Implement a feedback loop to learn from past decisions and adjust your reasoning process to avoid repeating mistakes. Reflect on your reasoning and consider feedback for continuous improvement in future evaluations. 12. **Keyword and Phrase Analysis**: Perform a detailed analysis of keywords and phrases. Compare these with the input question to justify classification decisions based on the presence or absence of domain-specific keywords. 13. **Consideration of Counterarguments**: Consider alternative perspectives or counterarguments that might suggest a question is out-of-domain. Address these counterarguments to provide a balanced and comprehensive analysis. ### Response Format: - **Decision**: [In-Domain/Out-of-Domain] - **Explanation**: [Provide a clear reason based on the question and knowledge base.]
@@
Figure 16: This prompt is designed to force GPT-4o to act an out-of-domain detector. Optimized by Yuksekgonul et al. (2025).
@@ Appendix C Further Results
C.1 Further Domains Results
@@
Method C19 SU Hist Cryp Chess Cook Astro Fit Anime Lit Bio Music Film Fin Law Comp Avg. Avg. #PCs
EVR criterion
$\epsilon$-ball 0.942 0.940 0.899 0.913 0.916 0.916 0.933 0.941 0.866 0.898 0.915 0.908 0.896 0.905 0.869 0.884 0.909 7
$\epsilon$-cube 0.937 0.928 0.894 0.912 0.906 0.910 0.928 0.933 0.850 0.901 0.908 0.906 0.896 0.898 0.863 0.869 0.902 6
$\epsilon$-rect 0.850 0.936 0.893 0.896 0.867 0.894 0.918 0.898 0.838 0.874 0.879 0.872 0.896 0.869 0.851 0.847 0.880 4
LogReg 0.981 0.967 0.963 0.966 0.959 0.983 0.975 0.967 0.939 0.953 0.930 0.927 0.940 0.938 0.903 0.917 0.951 140
SVM 0.985 0.972 0.962 0.967 0.960 0.982 0.974 0.967 0.941 0.946 0.929 0.933 0.938 0.940 0.922 0.921 0.952 143
GMM 0.937 0.964 0.928 0.976 0.948 0.977 0.977 0.950 0.878 0.905 0.904 0.892 0.923 0.880 0.877 0.872 0.924 40
p-values criterion
$\epsilon$-ball 0.985 0.944 0.933 0.908 0.923 0.922 0.922 0.936 0.844 0.913 0.870 0.912 0.887 0.884 0.849 0.852 0.901 13
$\epsilon$-cube 0.951 0.944 0.932 0.897 0.917 0.908 0.930 0.940 0.861 0.898 0.852 0.904 0.879 0.885 0.832 0.862 0.900 15
$\epsilon$-rect 0.942 0.904 0.775 0.772 0.855 0.778 0.834 0.767 0.815 0.888 0.714 0.864 0.848 0.803 0.780 0.845 0.824 5
LogReg 0.966 0.964 0.966 0.968 0.961 0.984 0.975 0.970 0.932 0.929 0.937 0.941 0.942 0.944 0.908 0.925 0.951 125
SVM 0.976 0.976 0.965 0.969 0.961 0.985 0.975 0.970 0.942 0.933 0.939 0.944 0.943 0.946 0.916 0.922 0.954 136
GMM 0.942 0.960 0.951 0.981 0.959 0.980 0.971 0.962 0.889 0.895 0.948 0.917 0.919 0.935 0.908 0.878 0.937 65
Baselines
Mahalanobis 0.990 0.800 0.947 0.976 0.961 0.981 0.975 0.957 0.920 0.845 0.945 0.891 0.945 0.929 0.924 0.895 0.930 768
kNN 0.976 0.976 0.955 0.982 0.975 0.977 0.979 0.963 0.917 0.919 0.943 0.945 0.952 0.940 0.909 0.924 0.952 768
ViM 0.937 0.808 0.844 0.920 0.896 0.929 0.929 0.909 0.759 0.803 0.742 0.882 0.887 0.798 0.778 0.792 0.851 768
ODIN 0.966 0.976 0.961 0.971 0.962 0.973 0.965 0.973 0.953 0.945 0.961 0.948 0.955 0.946 0.931 0.928 0.957 768
Table 7: Detailed OOD detection accuracy across 16 domains. Methods are grouped by feature-selection criterion (EVR vs p-values). The final columns report the average performance across domains and the average number of principal components retained.
Then, we show the detailed results for each combination of positive and negative samples in the training set and for different test sets across our three datasets: COVID-19 (always treated as positive), LLM-attack, and 4chan. The results are shown in Tables 8-15.
@@
For robustness, we experimented with different models as embedding extractors. More specifically, we employed ModernBERT with a pooling mechanism to extract sentence embeddings and NovaSearch/stella_en_400M_v5, which has shown state-of-the-art results in many NLP tasks. The experiments focused on the p-values criterion and the training sets of COVID-19 samples as positive and a mixture of all datasets as negative. The results are shown in Table 16,  17,  18 for the testing sets of a mixture of all datasets, 4chan, and LLM-Attack, correspondingly. As observed, the performance does not have significant fluctuations among the different models, indicating the effectiveness of a simplistic model that is used throughout our main findings.
@@
C.2 Further Relevance & Correctness Results
@@
[…]
@@
We conducted additional experiments focusing exclusively on Correctness to gain deeper insights into potential errors. Initially, we annotated all samples independently. Subsequently, we replaced the LLM-as-a-Judge with the Claude-3.5-Sonnet model. Interestingly, the new judge independently identified the necessity of a third category, "N/A", and classified 18 cases accordingly. To further validate this, we also allowed ourselves to mark "N/A" cases independently of Claude. This resulted in a high agreement between our annotations and the model’s judgments, indicating 12 common cases as "N/A". The results of both approaches are presented in Table 23.
@@
Table 23: Results for Correctness with samples annotated entirely by us and by Claude. "N/A" cases mean that these cases cannot be judged scientifically. 12 samples were noted as "N/A" from both our and Claude annotation.
As shown in the prompts in Fig.6 and Fig.7, two different prompts were used for the Generator component. This approach was necessary because using a single prompt proved ineffective in maximizing the number of questions that could be answered. To address this, we first employed the prompt that yielded the highest number of responses (Fig.6). The unanswered queries from this round were then processed using a second prompt (Fig.7). In the initial round, 281 out of 300 queries were successfully answered. The remaining 19 queries were reattempted using the second prompt across 10 iterations, resulting in 16 additional responses. Table 24 illustrates the results when all 300 samples are utilized. Table 25 presents the results when only the initial 281 answered queries are considered. Table 26 reports the results for the 16 additional responses generated by the second prompt, and Table 27 provides the final results for all 297 answered queries. The three unanswered queries were met with the response: "I can’t assist with this question."
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
In Table 30, we present the complete results for all developed methods evaluated on the 300 samples from our second study. This table complements Table 3, which exclusively reports the results for the GMM as it achieved the best performance. Similarly, in Table 31, we provide the results for the additional variations of the GPT-4o evaluator that we experimented with—specifically, the setting where only positive samples were included in the prompt and the 20-shot learning approach. Finally, in Table 32, we present our prompt optimization results. For optimization purposes, we developed 3 additional prompts (Fig. 13 - 15). The last two (Fig. 14 and  15) make use of zero-shot CoT (Chain-of-Thought) Kojima et al. (2022). For even further optimization, we utilize the methodology of Yuksekgonul et al. (2025), which results in the prompt of Fig. 16. As we observe, our best results are achieved by the prompt of the main paper. In addition, we leveraged this best prompt and developed an extensively utilized method. Specifically, we ran five passes of our best prompt using a temperature of 0.7. We then applied a majority vote strategy across the five predictions to reduce uncertainty in the classification of in-domain (ID) vs. out-of-domain (OOD) queries. The updated result is included as a new row (“UA” – Uncertainty-Aware). We find that the UA setup correctly classifies one additional query in total compared to the Main Prompt setup, indicating a minor improvement. This suggests that while incorporating stochasticity and ensembling can add value, our proposed methods remain highly competitive, especially given their interpretability and computational efficiency.
@@
[…]
@@ Appendix D Meaning Of PCs
This manual examination was followed by a verification step using GPT-4o, which confirmed the identified patterns. In Tables 34 and 35, we present the discovered patterns for the COVID-19 and SU domains, respectively. Each table includes: (1) the experimental condition in which the PC was selected, (2) a brief label describing the dominant pattern that the PC appears to capture, and (3) three example queries with the highest scores along that PC.